OpenStudy (anonymous):

what is equivalent to (sin x + 1)(sin x − 1)?

OpenStudy (anonymous):

(sin(x) + 1)(sin(x) - 1) = sin^2(x) - 1 sin^2(x) + cos^2(x) = 1 sin^2(x) - 1 + cos^2(x) = 0 Therefore: sin^2(x) - 1 = -cos^2(x)

OpenStudy (anonymous):

-cos2x?

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (raden):